OpenStudy (anonymous):

What is the distance between the points (-2,1) and (3,7)?

OpenStudy (anonymous):

Do you know the distance formula?

OpenStudy (anonymous):

no thats why im confused hahah

OpenStudy (anonymous):

Do you know pythagorean theorem? a^2+b^2=c^2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

what would that help me with doesn't that have to do with finding like a triangle?

OpenStudy (anonymous):

okay so lets tart with that. Pythagorean theorem tells us how to find the hypotenuse of a right triangle when we are given the 2 legs. In this case, you have 2 poitns (x,y)

OpenStudy (anonymous):

ok ;]

OpenStudy (anonymous):

you can find the difference in x (x2-x1) and (y2-y1) to find the two legs which will form a triangle with the distance (hypotenuse in this case)

OpenStudy (anonymous):

so 6 over 5

OpenStudy (anonymous):

distance formula is

OpenStudy (anonymous):

the possible answers that i have are square roots

OpenStudy (anonymous):

so your two legs are 6 and 5. Now go back to pythagorean theorem. The square of the two legs = the square of the hypotenuse. (leg1)^2 +(leg2)^2 = (hypotenuse)^2

OpenStudy (anonymous):

thankyou it all makes sense now WOW